Find a point-slope equation of the line having the given slope and containing the given point
m=-6;(4,8)

Answer by mananth(16949)

m=-6;(4,8)
The general equation is
(y-y1) = m(x-x1)
whwre m is the slope and (x1,y1) are the co ordinates of a point
(y-8)=-6(x-4)
y-8=-6x+24
y=-6x+32